OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

put n = number of horses, they have 4n legs. if there are n horses there are 74 - n humans and they have 2(74-n) legs. so total number of legs its \[4n+2(74-n)\] and it is also 196 so solve \[4n+2(74-n)=196\] to get your answer

OpenStudy (anonymous):

oh so the answer is 24 horses and 50 humans

OpenStudy (anonymous):

am i right???

OpenStudy (anonymous):

\[4n+144-2n=196\] \[2n=52\] \[n=26\] may be i made a mistake. i got 26 horses let me check again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no i made a mistake, you are right

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

50 humans gives 100 legs, 24 horses gives 96 legs, total legs is 196 and total heads is 74 you are right!

OpenStudy (anonymous):

i appreciate you giving me the equations.

OpenStudy (anonymous):

ha lol you seem very suprised

OpenStudy (anonymous):

i have to figure out my mistake. that is all

OpenStudy (anonymous):

ok well i think i can get the equation to show you what i did.

OpenStudy (anonymous):

oh i see it \[2\times 74=148\] and i wrote \[2\times 74=144\] that was my error

OpenStudy (anonymous):

oh ok